3x+5(2x-11)=2+9x: 3x+5(2x-11)=2+9x
Answer by orca(336)

Do you mean that solving the equation 3x+5(2x-11)=2+9x for x?
If so, see the following steps:
Removing the parenthesis,we have
3x + 10x - 55 = 2 + 9x
3x + 10x - 9x = 2 + 55
4x = 57
x = 57/2